OpenStudy (anonymous):

A swimmer runs horizontally off a diving board with a speed of 3.73 m/s and hits the water a horizontal distance of 1.56 m from the end of the board. (a) How high above the water was the diving board?

OpenStudy (anonymous):

range= velocity * time of descent. \[1.56=3.73\times \sqrt{2h/10}\] on solving the equation, you will get the height

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

whats in the square root? soory i cant really see it?

OpenStudy (anonymous):

2h/g

OpenStudy (anonymous):

where "h" is the height and "g" is the acceleration due to gravity

OpenStudy (stormfire1):

You can also get the same answer this way: \[\frac{1.56 m/s}{3.73 m/s}=0.418s\]\[d = Vi + \frac{1}{2} g t^2=0 m.s +\frac{1}{2}(10 m/s^2)(0.418s)^2\]

OpenStudy (stormfire1):

Oops...the top should be 1.56m.

OpenStudy (anonymous):

got it both works, thanks